OpenStudy (anonymous):

I'm confused I don't even know where to start with this question help! For the given quadratic equation convert into vertex form, find the vertex, and find the value for x = 6. Show your work. y = -2x2 + 2x +2

OpenStudy (anonymous):

Well first you need to complete the square, do you know how to do that?

OpenStudy (anonymous):

umm . I dont remember

OpenStudy (anonymous):

first group the x terms together so you would have y= (-2x^2 +2x) +2

OpenStudy (anonymous):

the take half of 2 in 2x and square and add it inside the parenthesis. so what do you get?

OpenStudy (anonymous):

square it*

OpenStudy (anonymous):

so if i half it then it would be 1 1/2^2?

OpenStudy (anonymous):

it would just be 1 because half of 2 is 1 and then 1^2 is just 1

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so then it breaks up into x^2+2x+2 ?

OpenStudy (anonymous):

no it would just be (-2x^2 + 2x +1) + 2

OpenStudy (anonymous):

because you aren't familiar with completing the square we can use the formula -b/2a or we can stick to completing the square it's up to you

OpenStudy (anonymous):

-b/2a is a little easier

OpenStudy (anonymous):

ok sure show me please

OpenStudy (anonymous):

okay so what represents b in the equation?

OpenStudy (anonymous):

-2x^2

OpenStudy (anonymous):

no, an example would be y=ax^2 + bx + c

OpenStudy (anonymous):

only focus on the leading coefficient

OpenStudy (anonymous):

well in this case not the leading coefficient but just the coefficients

OpenStudy (anonymous):

so would the leading coefficient be -2? because it starts first?

OpenStudy (anonymous):

sorry i told you the wrong thing that's my fault. it would just be 2

OpenStudy (anonymous):

or just 2? lol its ok thats what i was reading as well

OpenStudy (anonymous):

okay so now -2 would be a. so using -b/2a, what would you get?

OpenStudy (anonymous):

ok just so im understanding correctly when plugging it into the formula -b/2a it becomes 2/-2?

OpenStudy (anonymous):

then the answer to that is -1

OpenStudy (anonymous):

not quite. when you plug it in you would get -2/2(-2) do you understand how i got that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we keep the same thingn in the formula

OpenStudy (anonymous):

okay then when you simplify it you get 1/2

OpenStudy (anonymous):

my responses arent being posted

OpenStudy (anonymous):

hmm that's weird i see your responses though, maybe it's your computer?

OpenStudy (anonymous):

but now you have x=1/2. i should've wrote the full formula but it's x=-b/2a so that's how i found x for the vertex, now all we need to find is y

OpenStudy (nikato):